CH3CO2H(I) + 2O2(q) ------------ 2CO2(q) + 2H2O(I)
Determine the number of moles of each product and reactant that would be needed or produced for each situation specified.
CH3CO2H(I)
O2 given 3.7moles
CO2 given 5.2 moles
H2O given 0.75 moles
1
Expert's answer
2020-06-08T15:41:19-0400

According to equation the moles of O2, CO2 and H2O are equal, as their coefficients are equal, the moles of CH3CO2H are two times less then the moles of any other species.

1) given 3.7 moles O2, then moles of CO2 = 3.7 moles, moles of H2O = 3.7 moles, moles of CH3CO2H = 3.7/2= 1.85 moles.

2) given 5.2 moles CO2, then moles of O2= 5.2 moles, moles of H2O=5.2 moles, moles of CH3CO2H = 5.2/2= 2.6 moles.

3) given 0.75 moles H2O, then moles of O2 = 0.75 moles, moles of CO2 = 0.75 moles, moles og CH3CO2H = 0.75/2= 0.375 moles.